OpenStudy (anonymous):

Find the flux of F = x^3 i  + y^3 j  + z^3k through the closed surface bounding the solid region x^2 + y^2 ≤ 4, 0 ≤ z ≤ 7, oriented outward.

OpenStudy (anonymous):

@wio can you help me with this please, let me tell you what I did first

OpenStudy (anonymous):

I got divF=3x^2+3y^2+3z^2

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}\int\limits_{0}^{7}(r^2+z^2)rdzdr d \theta \]

OpenStudy (anonymous):

times 3 and I got 4620pi

OpenStudy (anonymous):

but I got it wrong

OpenStudy (anonymous):

@dan815 can you help me out :)

OpenStudy (freckles):

well have you integrate r^3+z^2r with w.r.t z?

OpenStudy (freckles):

you know starting with the inner most integral you have there

OpenStudy (anonymous):

@freckles where do you get r^3+z^2r from?

OpenStudy (freckles):

\[(r^2+z^2)r =r^3+z^2r \text{ by distributive property }\]

OpenStudy (anonymous):

how would I make green's theorem work here

OpenStudy (anonymous):

well yeah when I do the integral I get 4620 pi

OpenStudy (jhannybean):

yeah I forgot it's only for line integrals, whereas this one is a volume since z ranges from 0 to 7.

OpenStudy (anonymous):

i'm wondering if i did my limits of integration right

OpenStudy (anonymous):

@ganeshie8 can you help me out with this one please

ganeshie8 (ganeshie8):

try 1540pi

ganeshie8 (ganeshie8):

bounds and everything looks good, you made a mistake while evaluating the integral thats all. btw, at this point you should let wolfram do the dumb work.. important thing is setting up the integral correctly, not evaluating it manually. http://www.wolframalpha.com/input/?i=%5Cint%5Climits_%7B0%7D%5E%7B2%5Cpi%7D%5Cint%5Climits_%7B0%7D%5E%7B2%7D%5Cint%5Climits_%7B0%7D%5E%7B7%7D%283r%28r%5E2%2Bz%5E2%29%29+dzdrd%5Ctheta

OpenStudy (jhannybean):

Kind of just relearned something myself, haha. Finding that normal vector.

ganeshie8 (ganeshie8):

surface integral ? that would be really painful as we need to find 3 normals and work 3 integrals... but it would be a nice practice for sure :)

OpenStudy (anonymous):

but don't I need to multiply 1540 pi by 3 because of the divF = 3(x^2+y^2+z^3)

OpenStudy (anonymous):

z^2 not z^3

ganeshie8 (ganeshie8):

it was already multiplied

ganeshie8 (ganeshie8):

now i see how u got 4620, youhave multiplied 3 two times

OpenStudy (anonymous):

upps haha I multiplied it by 3 two times

OpenStudy (anonymous):

yeah haha

OpenStudy (anonymous):

silly mistake

OpenStudy (anonymous):

thanks for checking my work @ganeshie8

ganeshie8 (ganeshie8):

np:) are you done with the two problems that you run off from yesterday ?

OpenStudy (anonymous):

so I have another question why if the flux id out of the close surface and not through the closed surface

OpenStudy (anonymous):

no I'm not done @ganeshie8 I'm stuck with them

ganeshie8 (ganeshie8):

think of it as fluid flow

OpenStudy (anonymous):

would it be just negative because is out

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

it helps to think of the given vector field as "velocity field" in 3d when working with divergence

ganeshie8 (ganeshie8):

flux is same as the amount of fluid passing through the surface per unit time

OpenStudy (anonymous):

so the flux would be the same?

ganeshie8 (ganeshie8):